# CITADEL: A Semi-Supervised Active Learning Framework for Malware Detection Under Continuous Distribution Drift

### v1

@@ Abstract
Android malware evolves rapidly, leading to concept drift that degrades the performance of traditional machine learning (ML)-based detection systems. While recent approaches incorporate active learning and hierarchical contrastive loss to handle this drift, they remain fully supervised, computationally expensive, and perform poorly on real-world datasets with long temporal spans. In particular, our evaluation highlights these limitations, particularly on LAMDA, a 12-year longitudinal dataset exhibiting substantial distributional shifts. Moreover, manual expert labeling cannot scale with the daily emergence of over 450,000 new malware samples, leaving most samples unlabeled and underutilized.
To address these challenges, we propose CITADEL, a robust semi-supervised active learning framework for Android malware detection. To bridge the gap between image-domain semi-supervised learning and binary feature representations of malware, we introduce malware-specific augmentations, Bernoulli bit flips and masking, that simulate realistic drift behaviors. CITADEL further integrates supervised contrastive loss to improve boundary sample discrimination and combines it with a multi-criteria active learning strategy based on prediction confidence, -norm distance, and boundary uncertainty, enabling effective adaptation under limited labeling budgets. Extensive evaluation on four large-scale Android malware benchmarks—APIGraph, Chen-AZ, MaMaDroid, and LAMDA demonstrates that CITADEL outperforms prior work, achieving F1 score of over 1%, 3%, 7%, and 14% respectively, using only 40% labeled samples. Furthermore, CITADEL shows significant efficiency over prior work incurring faster training and fewer operations.
@@ I Introduction
The Android ecosystem, with over 3 billion active devices, has become a primary target for malware developers [1, 2]. The AV‑TEST Institute reports around 5 million new Android malware and potentially unwanted applications (PUAs) annually [3], posing a growing challenge to reliable threat detection. Traditional methods such as sandbox analysis and signature matching are labor-intensive, brittle to obfuscation, and struggle to scale [4, 5]. As a scalable alternative, machine learning (ML)-based detection systems have gained widespread adoption in both academia and industry [6, 7, 8], including integration by vendors such as Avast [9]. These systems learn from structural and behavioral features, and have demonstrated strong performance and generalization [10, 11].
@@
A core challenge in ML-based malware detection is concept drift–the change in the statistical properties of malware and benign samples over time. State-of-the-art concept drift adaptation method leverage active learning (AL) with hierarchical contrastive learning and uncertainty-based sample selection [11]. Though effective on short-span benchmarks—yielding 92.4% F1 on the 7-year APIGraph dataset [13], its performance drops sharply on longer temporal spans, dropping to 43% on the 12-year LAMDA benchmark [15] and further to 30% on a 12-year curated dataset derived from MaMaDroid [14] feature (see Figure 1).
@@
Furthermore, with over 450,000 new malware samples emerging daily [16], producing high-quality labels for these enormous volume is time, resource, and budget intensive; human analysts can review only around 80 samples per day [17], and vendors may take weeks to agree on newly discovered variants [18, 19]. Existing concept drift adaptation methods [20, 11] rely heavily on supervised learning, assuming continuous access to labeled data and focusing primarily on short-term evaluations. Such assumptions do not hold in realistic, long-term drift scenarios, where labeled data are scarce and the malware distribution evolves unpredictably. These challenges highlight the need for semi-supervised learning (SSL), which can leverage the abundance of unlabeled samples to maintain adaptability and performance in dynamic, label-scarce environments [21].
@@
To address long-term malware concept drift adaptation under limited labeling budgets, we propose CITADEL, a SSL framework adapted from FixMatch [22], specifically tailored for Android malware detection under distribution shift. Although prior work utilize FixMatch [12], their goal is to tackle label noise and data imbalance, which does not address the challenges of concept drift. Moreover, FixMatch is originally designed for computer vision domain and it suffers significantly to offer even reasonable performance for malware detection (see Table III). As such, we improve it for malware feature representations with domain-relevant transformations—Bernoulli bit flip and Bernoulli feature mask.
In practice, there is no deterministic pattern describing how feature distributions evolve over time [7], as new malware families may stop using certain API calls, permissions, or intent filters, while adopting others. As such, we hypothesize that Bernoulli-based transformations would be better fit to model the unpredictable nature of adversary induced distribution shifts in malware evolution. To simulate the uncertainty of adversary induced drift, we adopt a probabilistic approach using the Bernoulli distribution to randomly select features for modification. In Bernoulli bit flip, a randomly chosen subset of binary features (such as API calls or permissions) is flipped turning a previously used feature “off” or an unused one “on.” This simulates behavioral changes where malware starts or stops using specific functionalities. For Bernoulli feature mask deactivates a randomly selected subset of features, representing the disappearance of certain patterns (such as deprecated API calls or obsolete permissions) in newer malware variants. Since static features cannot capture entirely new APIs introduced in the future, this feature-space perturbation provides a principled way to simulate evolving distributions.
Our analysis show that CITADEL learns feature representations that effectively capture temporal drift. Adapting to such evolving patterns requires AL to selectively query the most informative samples [11]. Furthermore, malware drift is often abrupt and unpredictable, making such ordered progression ineffective. Many drifted samples tend to lie near the decision boundary—regions where the model is most uncertain and prone to misclassification [23, 24]. Additionally, outlier samples often appear as distant points in the latent space, further challenging accurate classification [20]. To address these issues, we propose a multi-criteria sample selection strategy that prioritizes: (i) low-margin samples (small confidence differences between top predictions), (ii) samples with high -norm distances, and (iii) low-confidence predictions. Together, these criteria target samples that are both ambiguous and influential for improving decision boundaries.
@@
Moreover, our analysis shows that boundary samples where benign and malware features overlap are particularly challenging to classify. To better handle such drifted and ambiguous cases, we introduce an auxiliary objective that promotes better separation of such ambiguous cases, complementing the semi-supervised loss. Figure 3 illustrates the overall architecture of CITADEL. Our extensive evaluation shows that CITADEL attains strong performance, achieving F1 scores of 93.5% on APIGraph, 82.7% on Chen-AZ, 44.9% on MaMaDroid, and 77.7% on LAMDA. In addition to accuracy, CITADEL provides substantial efficiency gains, offering up to 24 faster training and requiring 13 fewer operations than prior methods.
@@
We propose CITADEL, a SSL framework for Android malware detection under long-term concept drift. CITADEL introduces malware domain-specific augmentations, Bernoulli bit flip and Bernoulli feature mask.
@@
To better discriminate ambiguous boundary samples, we integrate a supervised contrastive loss [25] alongside the standard SSL objective, promoting clearer separation of low-margin samples from different classes in latent space.
@@
We propose a Multi-Criteria sample selection strategy for AL that selects highly informative samples by jointly considering boundary margin, -norm distance, and low-confidence predictions—enabling efficient adaptation under labeling budget constraints.
@@
We evaluate CITADEL on four Android malware benchmarks—APIGraph [13], Chen-AZ [11], MaMaDroid [14], and LAMDA [15]—under static and AL settings. CITADEL consistently outperforms prior methods, achieving F1 gains of on APIGraph, on Chen-AZ, on MaMaDroid, and on LAMDA. It also offers up to faster training and requires fewer operations, making it effective and efficient for deployment.
II Background and Related Work
In this section, we outline prior research on handling concept drift in malware analysis, including detection, adaptation, and learning strategies such as semi-supervised, active, curriculum, and continual learning.
Concept Drift Detection
Detecting concept drift is crucial for maintaining the long-term effectiveness of malware classifiers as threats evolve. Strategies in incremental learning [26, 27] aim to maintain accuracy while minimizing retraining. Key strategies include Gama et al. [28]’s drift detection methods to track model performance, Yang et al. [20]’s CADE framework using contrastive learning for dimensionality reduction and drift detection, and Barbero et al. [29]’s TRANSCENDENT framework that uses statistical analysis to quantify prediction credibility. OWAD [30] proposes an unsupervised method for detecting normality shift—a form of concept drift in zero-positive anomaly detection where only normal data is available. These methodologies enable reliable adaptation and decision-making without frequent retraining.
Concept Drift Adaptation
Upon detecting concept drift, effective adaptation is crucial for maintaining classifier performance against evolving malware and goodware (benign). Several prior works have jointly addressed both drift detection and adaptation. Jordaney et al. [31] developed a framework for early adaptation using kernel-based drift detection. Mishra et al. [32] used clustering to find feature shifts in malware. Chen-AL [11] applied hierarchical contrastive and active learning for efficient adaptation, lowering labeled data needs while improving F1 scores. Domain adaptation methods has also been studied for drift adaptation. Specifically, Pan et al. [26] started this, followed by Li et al. [27], who used graph-based techniques to focus on drift-invariant features resilient to obfuscation. These efforts underscore the importance of drift detection and adaptation for reliable malware classification. On the other hand, OWAD [30] selects key drift-inducing samples via an optimization-based explainer, reducing labeling cost. It updates the model using parameter-wise regularization that preserves important knowledge while learning new patterns.
Semi-Supervised Learning (SSL)
SSL has become increasingly valuable for malware detection, where labeled samples are limited but unlabeled data are abundant. Prior studies have shown that SSL can enhance detection accuracy while reducing manual effort. DeepReflect [33] integrates analyst feedback to minimize reverse engineering, while Kan et al. [34] demonstrate that pseudo-labeling helps classifiers adapt to new threats without frequent retraining. Apruzzese et al. [21] show that leveraging unlabeled data improves detection across diverse malware families. Collectively, these works emphasize SSL’s potential to build scalable and adaptive malware detection systems.
SSL has also gained traction in computer vision (CV) due to abundant unlabeled data and the high cost of manual annotation. Frameworks such as FixMatch [22], FlexMatch [35], FreeMatch [36], SoftMatch [37], and Dash [38] follow a common pipeline—generating pseudo-labels from weakly augmented inputs, filtering them by confidence, and enforcing consistency on strongly augmented variants. Each method refines this process through different confidence strategies, including class-wise, sample-wise, or momentum-based adjustments. However, image augmentations like flips or color shifts are inapplicable to binary malware features.
Active Learning in Malware Detection
Active learning (AL) is used in security to reduce labeling costs and keep models adaptable. Instead of labeling all data, AL queries uncertain samples for annotation, aiding rapid retraining with new malware. This is useful in malware detection, where expert labeling is costly and malware evolves quickly. Fang et al.’s BODMAS framework [39] found uncertainty sampling boosts malware detection performance. Zhou et al.’s APIGraph [40] enhanced this with meta-features to merge similar API calls, improving Android malware detection’s robustness with AL.
Other learning paradigms have been explored for malware classification. Casandra [41] uses online learning, updating the classifier with each newly labeled app for continual adaptation—though this is computationally demanding and impractical in resource-limited settings. DroidEvolver [42] applies self-training, assigning pseudo-labels to unlabeled data to reduce manual effort; however, Demetrio et al. [43] warn this can cause self-poisoning, where model errors are reinforced.
Curriculum Learning (CurL)
@@ II Background and Related Work
CurL primarily focuses on structuring training from easy to hard examples to improve model stability and convergence. Zhou et al. [44] adjusts sample difficulty based on model uncertainty, allowing the curriculum to evolve during training. Shashidhar et al. [45] present a model curriculum for teaching malware analysis, combining theoretical foundations with practical exercises such as reverse engineering and dynamic analysis which is out of scope in this paper.
Continual Learning (CL)
Unlike active learning, which aims to reduce annotation costs by selecting informative samples, CL focuses on maintaining performance across evolving tasks or data distributions by mitigating catastrophic forgetting [46]. Rahman et al.[47] first demonstrated that CL methods developed for computer vision (CV) perform poorly in the malware domain, highlighting the need for a domain specific CL framework. Recently, MalCL[48] proposes a GAN-based Class-IL framework that leverages feature matching loss and intermediate-layer-guided replay sample selection to improve generative replay and reduce forgetting. Furthermore, MADAR [49] proposes a diversity-aware replay strategy that balances representative and anomalous samples, capturing both inter- and intra-family variations to improve performance.
III Preliminaries
III-A Problem Setup and Notations
@@
We formulate our task as a semi-supervised malware detection problem under concept drift, where the data distribution of the malware (or benign) samples evolves over time. Let denote the input space of feature representations extracted from Android application (e.g., API call sequences or DREBIN features [6]), where 0 corresponds to absence and 1 corresponds to the presence of the feature of a sample; and let denote the label space, where corresponds to benign software and to malware.
Concept drift, arising from the evolving behavior of malware over time. We illustrate this in Figure 2 using the t-SNE projections latent space representations of various training samples of the APIGraph datasets that are spanned across several months. Specifically, some malware samples align closer to benign samples, and vice versa, leading to misclassifications of both classes. Additionally, certain samples deviate significantly from both the classes, forming outliers. Furthermore, boundary samples are particularly susceptible to shifts in data distribution and may challenge classifier robustness.
At each time step , the detector receives a small labeled dataset and a large pool of unlabeled data , where . The goal is to learn a malware detector with parameters that estimates the posterior probability under the current data distribution , such that the detection performance generalizes well even as shifts over time due to concept drift. Concept drift refers to a temporal shift in the joint distribution such that:
This change can be caused by new malware variants, evasion techniques, or benign software mimicking malicious behavior and vice versa. Drift can affect the input distribution (covariate drift), the conditional distribution (real concept drift), or both. To address evolving distributions, the learning framework must support: (1) adaptive updates to the model without full retraining, (2) robust use of unlabeled data under shifting domains, and (3) efficient selection of informative samples when labeling budgets are constrained.
III-B Motivating Example

### v2

@@ Abstract
Android malware detection systems suffer severe performance degradation over time due to concept drift caused by evolving malicious and benign app behaviors. Although recent methods leverage active learning and hierarchical contrastive loss to address drift, they remain fully supervised, computationally expensive, and ineffective on long-term real-world benchmark. Moreover, expert labeling does not scale to the monthly emergence of nearly 300K new Android malware samples, leaving most data unlabeled and underutilized.
To address these challenges, we propose CITADEL, a semi-supervised active learning framework for Android malware detection. Existing semi-supervised methods assume continuous and semantically meaningful input transformations, and fail to generalize well to high-dimensional binary malware features. We bridge this gap with malware-specific augmentations, Bernoulli bit flips and feature masking, that stochastically perturb feature to regularize learning under evolving malware distributions. CITADEL further incorporates supervised contrastive loss to improve boundary sample discrimination and combines it with a multi-criteria active learning strategy based on prediction confidence, -norm distance, and boundary uncertainty, enabling effective adaptation under constrained labeling budgets. Extensive evaluation on four large-scale Android malware benchmarks—APIGraph, Chen-AZ, MaMaDroid, and LAMDA, demonstrates that CITADEL outperforms prior work, achieving F1 score of over 1%, 3%, 7%, and 14% respectively, using only 40% labeled samples. Furthermore, CITADEL shows significant efficiency over prior work incurring faster training and fewer operations.
@@ 1 Introduction
The Android ecosystem, with over 3 billion active devices, has become a primary target for malware developers [samat2022android, abusnaina2025exposing]. The AV‑TEST Institute reports around 5 million new malware and potentially unwanted applications (PUAs) annually [avtest], posing a growing challenge to reliable threat detection. Traditional methods such as sandbox analysis and signature matching are labor-intensive, brittle to obfuscation, and struggle to scale [rastogi2013droidchameleon, zhou2012dissecting]. As a scalable alternative, machine learning (ML)-based detection systems have gained widespread adoption in both academia and industry [arp2014drebin, tesseract, domingues2021incremental], including integration by vendors such as Avast [avast2021ai]. These systems learn from structural and behavioral features, and have demonstrated strong performance and generalization [transcendingtranscend, chen2023continuous].
@@
A core challenge in ML-based malware detection is concept drift–the change in the statistical properties of malware and benign samples over time, which degrades detector performance. Prior work has addressed concept drift using contrastive learning, with active learning (AL) incorporated through uncertainty-based sample selection [chen2023continuous]. However, these approaches are evaluated under limited temporal settings. Although Chen-AL, as the state of the art, performs effectively on short temporal spans—achieving a 92.4% F1-score on the 7-year APIGraph dataset [api_graph_dataset]—its performance degrades substantially as the temporal span increases, dropping to 43% on the 12-year LAMDA benchmark [haque2025lamda] and further to 30% on a 12-year curated dataset derived from MaMaDroid [mamadroid] features (Figure 1).
@@
Furthermore, the scale of Android malware makes sustained labeling impractical. With over 300,000 new android malware samples emerging each month [avtest], producing high-quality labels is costly and slow; human analysts can review only around 80 samples per day [miller2016reviewer], and vendors may take weeks to agree on newly discovered variants [zhu2020measuring, kantchelian2015better]. Despite these constraints, existing concept drift adaptation methods [yang2021cade, chen2023continuous] rely heavily on supervised learning, implicitly assuming continuous access to labeled data and focusing on short-term evaluations. Such assumptions do not hold in realistic, long-term drift settings, where labels are scarce and the malware distribution evolves unpredictably. These challenges highlight the need for semi-supervised learning (SSL), which can leverage abundant unlabeled data to sustain detection performance under long-term distribution shifts [apruzzese2022sok].
@@
To address long-term malware concept drift adaptation under limited labeling budgets, we propose CITADEL, an SSL framework adapted from FixMatch [sohn2020fixmatch] and tailored for Android malware detection under distribution shift. While prior work applies FixMatch to malware [wu2023grim], it primarily tackles label noise and data imbalance rather than concept drift. Moreover, FixMatch is originally designed for computer vision and performs poorly when directly applied to malware feature representations (see Table 3). To address this gap, CITADEL adapts FixMatch with domain-specific feature transformations—Bernoulli bit flip and Bernoulli feature mask, designed for structured binary malware features.
In practice, malware evolution introduces adversary-driven changes in the feature space that are difficult to predict. Malware families may stop using certain API calls, permissions, or intent filters, while adopting others to realize the same behavior [api_graph_dataset]. Although such drift is often structured, to the best of our knowledge, prior works do not identify generalizable patterns that characterizes how static feature distributions evolve over time [mcfadden2025drmd, kaya2025ml, transcend, tesseract, yang2021cade]. As a result, we adopt a stochastic approach to modeling feature variation.
In particular, CITADEL employs Bernoulli-based [ross2014introduction] feature augmentations to inject controlled noise into the feature space. In Bernoulli bit flip, a randomly selected subset of binary features is flipped, simulating behavioral changes where malware starts or stops using specific functionalities. In Bernoulli feature mask, a randomly selected subset of features is deactivated, modeling the disappearance of observable patterns such as deprecated API calls or obsolete permissions in newer malware variants. When applied with low perturbation probabilities, these transformations introduce mild variation without significantly altering class semantics, whereas excessive perturbation can distort feature representations and degrade detection performance (see Appendix C). While static features cannot capture entirely new APIs introduced in the future [tesseract], these feature-space augmentations act as noise injection that regularizes the model against feature instability, encouraging robustness to unpredictable distribution shift.
While feature-space augmentations improve robustness to evolving malware behavior, effective adaptation also depends on which samples are selected for labeling over time. Adapting to such evolving patterns therefore requires active learning (AL) to selectively query informative samples [chen2023continuous]. However, malware drift is often abrupt and unpredictable, limiting the effectiveness of selection strategies that assume gradual or ordered change. Under these conditions, many drifted samples lie near the decision boundary, where the model is most uncertain and prone to misclassification [ducoffe2018adversarial, vzliobaite2011active]. In addition, outlier samples often appear far from existing cluster in the latent space, further challenging accurate classification [yang2021cade].
To address these challenges, we propose a multi-criteria sample selection strategy that combines: (i) low-margin samples with small confidence gaps, (ii) samples with large -norm distances indicating distributional deviation, and (iii) low-confidence predictions. These criteria focus labeling effort on samples that are both ambiguous and influential for refining the decision boundary under distribution shift.
@@
Moreover, our analysis shows that samples near the decision boundary where benign and malware features overlap are particularly difficult to classify. To better handle such drifted and ambiguous cases, CITADEL incorporates an auxiliary objective that promotes improved class separation, complementing the semi-supervised loss. Figure 3 provides an overview of the architecture of CITADEL. Our extensive evaluation shows that CITADEL achieves strong performance across multiple datasets, with F1 scores of 93.5% on APIGraph, 82.7% on Chen-AZ, 44.9% on MaMaDroid, and 77.7% on LAMDA. In addition, CITADEL improves training efficiency, achieving up to 24 faster training and requiring 13 fewer operations compared to prior methods.
@@
We propose CITADEL, a semi-supervised active learning (SSAL) framework for Android malware detection under long-term concept drift. CITADEL introduces domain-specific feature augmentations, Bernoulli bit flip and Bernoulli feature mask, for structured malware features.
@@
To better discriminate ambiguous boundary samples, we integrate a supervised contrastive loss [khosla2020supervised] with the standard SSL objective, improving class separation of low-margin samples in the latent space.
@@
We design a Multi-Criteria AL strategy that selects informative samples by jointly considering boundary margin, -norm distance, and low-confidence predictions—enabling efficient adaptation under labeling budget constraints.
@@
We evaluate CITADEL on four Android malware benchmarks—APIGraph [api_graph_dataset], Chen-AZ [chen2023continuous], MaMaDroid [mamadroid], and LAMDA [haque2025lamda]—under both static and AL settings. CITADEL consistently outperforms prior methods, achieving F1 gains of on APIGraph, on Chen-AZ, on MaMaDroid, and on LAMDA, while providing up to faster training and requiring fewer operations.
We construct a new dataset based on MaMaDroid [mamadroid] features spanning 12 years, mirroring the temporal characteristics of LAMDA [haque2025lamda].
2 Related Work
Concept Drift Detection and Adaptation. Concept drift is commonly detected with respect to the performance degradation of a trained model on future unseen data [tesseract, anoshift]. In malware detection, identifying concept drift is critical, as both malware and benign applications continuously evolve, leading to shifts in feature distributions. Prior work explored various strategies for concept drift adaptation. Several approaches leverage contrastive learning to identify distributional changes or out-of-distribution samples [chen2023continuous, yang2021cade, he2025combating]. For instance, CADE [yang2021cade] proposes a contrastive framework for detecting out-of-distribution malware samples, while Chen-AL [chen2023continuous] utilizes active learning with uncertainty-based sample selection to efficiently query labels from malware analysts under a limited budget. TRANSCENDENT [transcendent] employs conformal rejection to identify unreliable predictions and retrain the model to mitigate performance degradation. However, most of these methods lack evaluation under long-term, large-scale, and highly pronounced drift scenarios, limiting their applicability to real-world malware evolution.
Semi-Supervised Learning (SSL). SSL techniques are significantly valuable where labeled data are scarce. Recent SSL techniques are widely recognized for their outstanding performance, such as FixMatch [sohn2020fixmatch], FlexMatch [zhang2021flexmatch], FreeMatch [wang2022freematch], SoftMatch [chen2023softmatch], and Dash [xu2023dash], in computer vision. [wu2023grim] is the first to utilize SSL technique like FixMatch for malware detection on noisy label dataset but they did not address concept drift. DeepReflect [downing2021deepreflect] integrates analyst feedback to minimize reverse engineering, while Kan et al. [kan2021investigating] demonstrate that pseudo-labeling helps classifiers adapt to new threats without frequent retraining. Apruzzese et al. [apruzzese2022sok] show that leveraging unlabeled data improves detection across diverse malware families.
Active Learning (AL) in Malware Detection. AL reduces labeling costs while enabling malware detectors to adapt to evolving threats by querying only the most uncertain samples for expert annotation. This is particularly effective in malware detection, where labeling is expensive and malware behavior changes rapidly. Prior work, such as Chen-AL [chen2023continuous], shows that uncertainty sampling improves detection performance, while APIGraph [zhang2020enhancing] enhances robustness by merging semantically similar API calls. Other paradigms have also been explored such as Casandra [demontis2017casandra] applies online learning for continual adaptation but incurs high computational overhead, while DroidEvolver [nguyen2019droidevolver] uses self-training to reduce labeling effort at the risk of self-poisoning [demetrio2021explaining].
@@ 2 Related Work
Curriculum Learning (CurL). CurL primarily focuses on structuring training from easy to hard examples to improve model stability and convergence. Zhou et al. [zhou2020curriculum] adjusts sample difficulty based on model uncertainty, allowing the curriculum to evolve during training. Prior work proposed CurL for malware analysis [shashidhar2016teaching], combining theoretical foundations with practical exercises such as reverse engineering and dynamic analysis which is out of scope in this paper.
Continual Learning (CL). Unlike AL, which reduces annotation cost by selecting informative samples, continual learning (CL) aims to maintain performance under evolving data distributions by mitigating catastrophic forgetting [parisi2019continual]. Rahman et al. [continual-learning-malware] showed that CL methods developed for computer vision perform poorly in the malware domain, motivating domain-specific approaches, while MADAR [madar] explores replay-based strategies for malware evolution. However, CL addresses a different problem setting and is out of the scope of this work.
3 Preliminaries
3.1 Problem Setup and Notations
@@
We formulate our task as a semi-supervised malware detection problem under concept drift, where the distribution of the malware and benign samples evolves over time. Let denote the input space of feature representations extracted from Android applications; and let denote the label space, where 0 corresponds to benign software and 1 to malware. At each time step , the detector receives a small labeled dataset and a large pool of unlabeled data , where . The goal is to learn a malware detector with parameters that estimates a score corresponding to under the current data distribution , while maintaining strong detection performance as changes over time.
Concept drift arises from the emergence of new malware variants, evolving evasion techniques, and benign software mimicking malicious behavior. Such drift can affect the input distribution (covariate drift), the conditional distribution (real concept drift), or both. We model concept drift as a temporal changes in the joint distribution such that:
Figure 2 illustrates this phenomena using 2d projections of latent space representations from APIGraph samples, spanned across several months. The visualization depicts increased overlap between malware (red) and benign (green) samples, along with our preference for outliers and boundary samples that are more sensitive to distribution shifts. In particular, boundary samples that are misclassified by the detector are captured using the softmax margin and are highlighted with black circles. The visualization is obtained by projecting penultimate-layer MLP embeddings into two dimensions using t-SNE with a perplexity of 30 and a fixed random seed (random_state = 42) for reproducibility.
We argue that addressing evolving distributions requires a learning framework that supports: (1) updating the model without full retraining, (2) effective use of unlabeled data as domains shift, and (3) efficient selection of informative samples when labeling budgets are constrained.
3.2 Motivating Example